OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 2x - 9y = 1 x - 4y = 1

OpenStudy (zzr0ck3r):

x = 1+4y 2(1+4y)-9y = 1 2+8y -9y = 1 -y=-1 y=1 so x = 1+4 x = 5